OpenStudy (anonymous):

Which is the best equation to use to solve this problem? The perimeter of a rectangle is 68 ft. The ratio of the length to the width is 9 : 8. What are the dimensions of the rectangle? A.2x+2x=68 B.9x+8x=68 C.9x/8x=68 D.9x+9x+8x+8x=68

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

length : width = 9:8 so let length = 9x, width = 8x ok?

OpenStudy (anonymous):

yes

hartnn (hartnn):

perimeter of rectangle = sum of sides = 2* length + 2* width

OpenStudy (anonymous):

l*w?

hartnn (hartnn):

length = l = 9x width = w = 8x perimeter = 2*(9x) + 2*(8x) = 68

hartnn (hartnn):

or 9x + 9x+ 8x+8x = 68

OpenStudy (anonymous):

ohhhhh so it is D??

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

yeahh thanks. :D

hartnn (hartnn):

welcome ^_^